8, 9, 10 PORTFOLIO ANALYSIS:

YEAR

ASSET HOTDOGS

BURGERS

ASSET WINGS

2016

3%

7%

14%

2017

7%

9%

13%

2018

10%

12%

12%

2019

12%

16%

11%

2020

13%

15%

10%

ALTERNATIVE

INVESTMENT

8

60% OF ASSET BURGERS AND 40% OF ASSET WINGS

9

55% OF ASSET HOTDOGS AND 45% OF ASSET BURGERS

10

33% OF ASSET HOTDOGS, 33% OF ASSET BURGERS AND 34% OF ASSET WINGS

PLEASE CALCULATE THE EXPECTED RETURN AND STANDARD DEVIATION FOR EACH OF THE THREE PORTFOLIO ALTERNATIVES.
Answer #1

In the questin it has given, 5 years return of 3 stocks namely Asset hotdogs, Burgers, Asset Wings.

Hence we have to calculate their Average Return on each portfolio..

Asset Hotdogs- (3+7+10+12+13)/5 = 45/5 = 9%.

Asset Burgers- (7+9+12+16+15)/5 = 59/5 = 11.8%

Asset Wings- (14+13+12+11+10)/5 = 60/5 = 12%

Caluculation of Average Deviation on each portfolio-

Formula = {(Return1- Average Return)2 + (Return2- Average Return)2........]1/2

Asset Hotdogs= [(3-9)2+(7-9)2+(10-9)2+(12-9)2+(13-9)2/5]1/2 = [13.2%]1/2 = 3.63%

Asset Burgers- [(7-11.8)2+(9-11.8)2+(12-11.8)2+(16-11.8)2+(15-11.8)2/5]1/2 = [11.76%]1/2 = 3.42%

Asset Wings- [(14-12)2+(13-12)2+(12-12)2+(11-12)2+(10-12)2/5]1/2 = [2%]1/2 = 1.41%

Formula-

Expected Return E(R) = W1*R1+W2*R2+ ...........+Wn*Rn

Alternative-8

Given, Weighted to Asset Burger- 60% and to Asset Wings- 40%

Expected Return =(Weight of Asset Burger * Return of Asset Burger) + ​(Weight of Asset Wings * Return of Asset Wings)

= (60%*11.8%) + (40%*12%) = 7.08%+4.80%= 11.88%

SD= (Weight of Asset Burger * SD of Asset Burger) + ​(Weight of Asset Wings * SD of Asset Wings)

= (60%*3.42%) + (40%*1.41%) = 2.052% + .564% = 2.616%

Alternative-9

Given, Weighted to Asset Burger- 45% and to Asset Hotdogs- 55%

Expected Return =(Weight of Asset Burger * Return of Asset Burger) + ​(Weight of Asset Hotdogs * Return of Asset hotdogs)

=(45%*11.8%)+(55%*9%) = 5.31%+4.95% = 10.26%

SD = (Weight of Asset Burger * SD of Asset Burger) + ​(Weight of Asset Hotdogs * SD of Asset hotdogs)

= (45%*3.42%)+(55%*3.63%) = 1.539% + 1.9965% = 3.536%

Alternative-10

Given, Weighted to Asset Hotdogs- 33%, Asset Burger- 33 % and to Asset Wings- 34%

Expected Return = ​(Weight of Asset Hotdogs * Return of Asset hotdogs) + (Weight of Asset Burger * Return of Asset Burger) + (Weight of Asset Wings * Return of Asset Wings)

= (33%*9%)+(33*11.8%)+(34%*12) = 2.97%+ 3.894% +4.08% = 10.944%

SD= (Weight of Asset Hotdogs * SD of Asset hotdogs) + (Weight of Asset Burger * SD of Asset Burger) + (Weight of Asset Wings * SD of Asset Wings)

= (33%*3.63%)+(33*3.42%)+(34%*1.41) = 1.198% + 1.123% + .480% = 2.801